# FGBERT: Function-Driven Pre-trained Gene Language Model for Metagenomics

### v1

@@ Abstract
Metagenomic data, comprising mixed multi-species genomes, are prevalent in diverse environments like oceans and soils, significantly impacting human health and ecological functions. However, current research relies on K-mer representations, limiting the capture of structurally relevant gene contexts. To address these limitations and further our understanding of complex relationships between metagenomic sequences and their functions, we introduce a protein-based gene representation as a context-aware and structure-relevant tokenizer. Our approach includes Masked Gene Modeling (MGM) for gene group-level pre-training, providing insights into inter-gene contextual information, and Triple Enhanced Metagenomic Contrastive Learning (TEM-CL) for gene-level pre-training to model gene sequence-function relationships. MGM and TEM-CL constitute our novel metagenomic language model FGBERT, pre-trained on 100 million metagenomic sequences. We demonstrate the superiority of our proposed FGBERT on eight datasets.
@@ 1 Introduction
Metagenomics, the study of mixed genomes of microbial communities in the environment (e.g. gut microbiomes or soil ecosystems) [37, 39], has revealed the critical role in fundamental biological processes like enzyme synthesis, gene expression regulation, and immune function [48, 36]. In light of this, deciphering the complex sequences of multiple species in metagenomics is vital for unraveling life’s mechanisms and advancing biotechnology [4, 33, 34].
Figure 1: Two types of complex relationships between gene sequences and functions in metagenomics. One-to-Many means that the same gene may display different functions based on the genomic context; for example, ATP synthase works differently in plants, heterotrophic bacteria, and humans. Many-to-One shows that multiple genes may perform the same function; for instance, different genes from different bacteria, e.g., Cpf1, Cas1, etc. produce the same resistance function within the immune system CRISPR.
In metagenomic data analysis, various computational methods have emerged. Traditional alignment-based methods rely on homology search against reference genome databases [53]. Machine learning-based methods [20, 31, 42] generally take K-mer as the basic processing unit and adopt its frequency or embedding as the input feature, which proved to be efficient alternatives to traditional methods. Recently, Transformer [57] has been widely used for modeling sequence data due to its ability to capture long-range dependencies. Several pre-trained genomic models [63, 10] utilize Transformer architectures to effectively model complex DNA contexts. However, these models, primarily designed for single-species analysis, face challenges in mixed multi-species metagenomic data. MetaTransformer [56] and ViBE [18] use K-mer embeddings, cater to specific metagenomic applications, with the former using Transformer for read classification and the latter pre-trained BERT to predict virus classes.
Yet existing genome or metagenomic pre-trained models still face many challenges when dealing with metagenomic data. First, how to encode a gene sequence with biological meaning? Typical pre-processing of genomic data often relies on K-mer to represent a local feature of a gene sequence. However, this approach, with its fixed and limited ‘vocabulary’ of K-mer combinations [63], falls short in capturing informative and global gene representations. Secondly, how to clarify the complicated relationships between gene sequences and their functions in metagenomic analysis? This challenge can be categorized as One-to-Many and Many-to-One problems, as opposed to single-species genome analyses (e.g., identifying regulatory elements in a particular genome). One-to-Many problem indicates that a single gene can exhibit various functions in different genomic contexts, underscoring the significance of inter-gene interactions in function regulation [59]. Figure 1a demonstrates that the ATP synthase exhibits specific functionalities in diverse organisms such as bacteria, plants, and humans. Conversely, Many-to-One problem implies that different genes can share the same function, emphasizing expression commonality [3]. As shown in Figure 1b, CRISPR [23] is an immune mechanism where different proteins such as Cpf1, Cas1, and Cas13, etc. serve this function.
To address these challenges, we propose FGBERT, a novel metagenomic pre-trained model equipped with the ability to encode contextually-aware and functionally relevant representations of metagenomic sequences. We propose a protein-based gene representation as a context-aware tokenizer to encode each gene sequence, which allows for a flexible token vocabulary for longer metagenomic sequences. This strategy leverages the inherent protein functional and structural information encoded in metagenomic data [48]; and, due to the redundancy of codon-to-amino acid mappings [12], since we adopt the embedding of protein sequences, the function remains unchanged despite potential mutations. With the new tokenizer, we propose a gene group-level pre-training task, named Masked Gene modeling (MGM), to mask some genes randomly and then pre-train a Roberta-like model to predict them, enhancing the understanding of inter-/intra-gene contextual relationships and how the same gene act in different genome contexts in One-to-Many scenarios. Furthermore, we propose Triplet Enhanced Metagenomic Contrastive Learning (TEM-CL) to identify genes with similar functions together with data augmentation and negative sampling module, boosting the model’s ability to recognize commonalities among gene functions in Many-to-One scenarios. MGM and TEM-CL constitute a joint pre-training model, which advances the co-representation learning of metagenomic gene sequences and functions.
@@
Contributions. To the best of our knowledge, this is the first metagenomic pre-trained model encoding context-aware and function-relevant representations of metagenomic sequences. To summarize: (1) We introduce a new idea of protein-based gene representations to learn biologically meaningful tokenization of long sequences. (2) Hinged on this idea, we propose MGM to model inter-gene relationships and TEM-CL to learn complex relationships between gene sequences and functions. (3) We conduct extensive experiments on four levels of downstream tasks. Our proposed FGBERT obtains state-of-the-art performance in comparison to a comprehensive suite of baselines.
@@ 2 Related Works
Research on Metagenomics Traditional alignment-based methods like MetaPhlAn5 [53], aim to match similarities between query sequences and known reference genomes and are common for taxonomic profiling. Advancements in deep learning have led to new methods like CNN-MGP [2] and DeepVirFinder [49], which use CNNs for gene and viral classifications with one-hot encoding. K-mer tokenization [14], employed in approaches like MDL4Microbiome [30], is a standard for DNA sequence characterization. Additionally, Virtifier [41] maps a nucleotide sequence using a codon dictionary, combined with LSTM to predict viral genes. DeepMicrobes [31] employs a self-attention mechanism, while DeepTE [58] uses K-mer inputs with CNNs for element classification, and Genomic-nlp [42] applies word2vec for gene function analysis. MetaTransformer [56] uses K-mer embedding for species classification with Transformer. For pre-training models, LookingGlass [20] uses a three-layer LSTM model for functional prediction in short DNA reads. ViBE [18] employs a K-mer token-based BERT model pre-trained with Masked Language Modeling (MLM) for virus identification.
Pre-Training on Genomics The BERT model, effective in DNA sequence characterization, is limited by the Transformer architecture’s computational burden. LOGO [60] addresses this by cutting off long sequences into 1-2kb sub-sequences. Enformer [6] combines extended convolution with Transformers for long human genomic data. GenSLMs [64] introduce hierarchical language models for whole-genome modeling. DNABERT [24], the first pre-trained model on the human genome that focuses on extracting efficient feature representations from gene sequences. DNABERT2 [63], its successor, uses Byte Pair Encoding on multi-species genomic data. NT [10] is trained on nucleotide sequences from human and other species genomes and evaluated on 18 genome prediction tasks. HyenaDNA [43] presents a long-range genome sequence model based on single-nucleotide polymorphisms on the human reference genome.
@@ 3 Methods
In this section, we detail the proposed pre-training framework FGBERT, which contains MGM and TEM-CL illustrated in Figure 2.
@@
Notation Given a metagenomic long sequence $\mathcal{X}$, after ultrasonic fragmentation and assembly techniques, N genes $\left\{{x_{i}}\right\}_{i=1}^{N}$ are obtained, which is tokenized to $g_{i}\in\mathbb{R}^{d}$ together with its corresponding label $y_{i}\in\mathbb{R}^{100}$, where $d$ is the token dimension. After that, we concatenate ${\left\{g_{i}\right\}}_{i=1}^{n}$ into a gene group $\mathcal{G}$. During the pre-training stage, each gene group $\mathcal{G}$ is passed to the context-aware genome language encoder $\mathcal{F}(\cdot)$ to obtain knowledge representations $z_{i}$ for the input token $g_{i}$, where $z_{i}=\mathcal{F}(g_{i})$. Then we apply a fully connected layer to obtain $\hat{z}_{i}$ from $z_{i}$, where $\hat{z}_{i}=\mathcal{H}(z_{i})$, $\mathcal{H(\cdot)}$ is a multi-classification head. Additionally, we integrate contrastive learning. For gene $x_{i}$, we introduce a data augmentation module to construct positive samples $x_{j(i)}$ and employ hard negative sampling strategy mining for negative sample construction $x_{k(i)}$.
3.1 Masked Gene Modeling (MGM)
Context-aware Tokenizer To address the first challenge, we propose the protein-based gene representation as a context-aware tokenizer, utilizing the ESM-2 [32] with 15 billion parameters. Naturally, this could be replaced by any Protein Language Model (PLM). This process starts with transforming the DNA sequence $x_{i}$ from metagenomic sequences $\mathcal{X}$ into the protein sequence by ENA software [17]. The AA sequence is then transformed into a 1280D normalized ESM-2 representation. An additional dimension is appended to encode gene orientation, resulting in 1281D gene representation $g_{i}$. These representations are concatenated to form gene groups $\mathcal{G}$ as shown in Figure 3. The conversion from DNA to AA sequences is motivated by the equivalence of certain degenerate DNA codes (i.e. non-‘ATCG’ symbols in $\mathcal{X}$) when translated into AA, underscoring the rationale for transitioning from DNA to AA sequences for clarity and consistency in representation. The gene-protein relationship inherited in metagenomic data itself [48] could be revealed by ESM-2’s comprehensive training data and protein structural and functional information, and thus ESM-2 representation preserves important intra-gene contextually-aware information. This approach bypasses traditional vocabulary limits and avoids raw sequence processing, this provides the flexibility to analyze longer metagenomic sequences.
Figure 3: Framework of our proposed Context-Aware Tokenizer.
Masked Gene Modeling With this new tokenizer, we propose MGM to pre-train the Gene Encoder. During pre-training, each token (gene) is masked with a 15% probability and predicted based on its genome context. The loss $\mathcal{L}_{MLM}$ is formulated as a cross-entropy loss:
$\mathcal{L}_{\text{MLM}}=\mathbb{E}_{g\sim\mathcal{G}}\mathbb{E}_{M}\sum_{i\in M}-\log p(g_{i}|{\mathcal{G}_{/M}})$ (1)
for any given gene representation $g$, a random set of indices, $M$ means the masked gene’s index set, the unmasked gene embeddings ${\mathcal{G}_{/M}}$ as the context.
In this way, BERT can learn inter-gene relationships within the genome, contributing to our understanding of gene functionality in diverse genomic contexts in One-to-Many scenarios. BERT implicitly learns intra-gene information encapsulated within the PLM representations, further enriching the model’s comprehension of gene characteristics.
What’s more, acknowledging gene polymorphism [47, 62], MGM emphasizes identifying multiple genes that may coexist at a single genomic site, denoted as $\hat{z}_{i}=[\hat{z}_{1},\hat{z}_{2},\hat{z}_{3},\hat{z}_{4}]$. This necessitates its ability to predict not the individual genes but also various gene combinations within the same site. Hence, two loss functions are combined into $\mathcal{L}_{\text{MGM}}$:
@@ 3.1 Masked Gene Modeling (MGM)
$\mathcal{L}_{\text{MGM}}=\frac{1}{N}\textstyle\sum\limits_{i=1}^{N}{(1-\frac{y_{i}^{T}\hat{z}_{i}}{\left\|y_{i}\right\|\cdot\left\|\hat{z}_{i}\right\|})}^{\gamma}+\\ \frac{\alpha}{N}\textstyle\sum\limits_{i=1}^{N}||\tilde{z}_{i}-\tilde{y}_{i}||_{2}^{2}$ (2)
@@
where $\gamma$ is a reconstruction loss with the scaled cosine error, and $\alpha$ is a weighting factor to balance the importance of the two loss functions. The first item is Feature Reconstruction Loss (FRL), which measures the distance between the model prediction $\hat{z}_{i}$ and its corresponding label $y_{i}$. The second item, Probability Prediction Loss (PPL), assesses the discrepancy between the embedding probability $\tilde{z}_{i}=\frac{e^{\hat{z}_{i}}}{\sum_{j=1}^{C}e^{\hat{z}_{j}}}$ and the true category probability $\tilde{y}_{i}=\frac{e^{y_{i}}}{\sum_{j=1}^{C}e^{y_{j}}}$ processed by the softmax function with C denoting the number of gene combinations and is set to 4.
3.2 Triplet Enhanced Metagenomic Contrastive Learning (TEM-CL)

### v2

@@ Abstract
Metagenomic data, comprising mixed multi-species genomes, are prevalent in diverse environments like oceans and soils, significantly impacting human health and ecological functions. However, current research relies on K-mer, which limits the capture of structurally and functionally relevant gene contexts. Moreover, these approaches struggle with encoding biologically meaningful genes and fail to address the One-to-Many and Many-to-One relationships inherent in metagenomic data. To overcome these challenges, we introduce FGBERT, a novel metagenomic pre-trained model that employs a protein-based gene representation as a context-aware and structure-relevant tokenizer. FGBERT incorporates Masked Gene Modeling (MGM) to enhance the understanding of inter-gene contextual relationships and Triplet Enhanced Metagenomic Contrastive Learning (TMC) to elucidate gene sequence-function relationships. Pre-trained on over 100 million metagenomic sequences, FGBERT demonstrates superior performance on metagenomic datasets at four levels, spanning gene, functional, bacterial, and environmental levels and ranging from 1k to 213k input sequences. Case studies of ATP Synthase and Gene Operons highlight FGBERT’s capability for functional recognition and its biological relevance in metagenomic research.
@@ 1 Introduction
Metagenomics, the study of mixed genomes of microbial communities in the environment (e.g. gut microbiomes or soil ecosystems) [41, 43], has revealed the critical role in fundamental biological processes like enzyme synthesis, gene expression regulation, and immune function [53]. This deepened understanding highlights the need to accurately interpret the intricate genetic information contained within these diverse communities. Consequently, deciphering the complex sequences of multiple species in metagenomics is vital for unravelling life’s mechanisms and advancing biotechnology [4, 37, 38].
Unlike traditional genomics focused on single species, metagenomics involves genetic material directly from environmental samples, posing significant challenges due to sample diversity and species abundance [40]. As shown in Fig. 1, the typical challenges in metagenomics are the presence of One-to-Many (OTM) and Many-to-One (MTO) problems. The OTM problem indicates that a single gene can exhibit various functions in different genomic contexts, underscoring the significance of inter-gene interactions in function regulation [63]. For example, ATP synthase displays distinct functionalities in diverse organisms such as bacteria, plants, and humans (Fig. 1a). Conversely, the MTO problem implies that different genes can share the same function, emphasizing expression commonality [3]. For example, the CRISPR immune mechanism involves various proteins like Cpf1, Cas1, and Cas13, each contributing to the same defensive function (Fig. 1b) [63, 3, 23].
Recently, various computational methods have emerged for genomic and metagenomic data analysis. However, these methods still face challenges when analyzing metagenomic data. Firstly, the Semantic Tokenizer problem. Most machine learning-based methods [20, 35, 47], taking K-mer counts, frequencies or embeddings as input features, providing efficient alternatives to traditional homology searches against reference genome databases. However, K-mer features often have limited representation ability and fail to capture global information. Secondly, the Function-Driven Modeling problem. Although recent Transformer models excel in modeling complex DNA contexts through long-range dependencies, they are predominantly tailored for single-species genomic analysis [61, 67, 10] and do not adequately address OTM and MTO challenges. This limitation impedes their ability to accurately model the intricate relationships between genes, their function across diverse genomic environments, and their connections among sequences with similar functions. Thirdly, the Low Generalization problem. Models like MetaTransformer [60] and ViBE [18], designed for specific tasks such as read classification and virus category prediction, fail to grasp the broader biological complexities of metagenomic data, limiting their generalization across diverse metagenomic tasks.
To address these challenges, we propose FGBERT, a novel metagenomic pre-trained model designed to encode contextually-aware and functionally relevant representations of metagenomic sequences. First, to solve the problem of encoding gene sequences with biological meaning, we propose a protein-based gene representation as a context-aware tokenizer, allowing for a flexible token vocabulary for longer metagenomic sequences. This strategy leverages the inherent protein functional and structural information encoded in metagenomic data [53], overcoming the limitations of K-mer methods and maintaining functional consistency despite potential mutations [12]. Second, we propose two pre-training tasks for function-driven modeling: Masked Gene Modeling (MGM) and Triplet Enhanced Metagenomic Contrastive Learning (TMC) to enhance the co-representation learning of metagenomic gene sequences and functions. Thirdly, FGBERT is pre-trained on over 100 million sequences, showcasing robust performance across diverse datasets spanning gene, functional, bacterial, and environmental levels.
@@
Contributions. In this work, we identify three key challenges in metagenomic analysis. To address these issues, we propose FGBERT. To the best of our knowledge, this is the first metagenomic pre-trained model encoding context-aware and function-relevant representations of metagenomic sequences. To summarize: (1) We introduce a new idea of protein-based gene representations to learn biologically meaningful tokenization of long sequences. (2) We propose MGM to model inter-gene relationships and TMC to learn complex relationships between gene sequences and functions. (3) We conduct extensive experiments across various downstream tasks, spanning gene, functional, bacterial, and environmental levels with input sizes from 1k to 213k sequences. FGBERT achieves SOTA performance.
Figure 1: Motivaion. Two types of complex relationships between gene sequences and functions in metagenomics. One-to-Many problem means that the same gene may display different functions based on the genomic context; for example, ATP synthase works differently in plants, heterotrophic bacteria, and humans. Many-to-One problem shows that multiple genes may perform the same function; for instance, different genes from different bacteria, e.g., Cpf1, Cas1, etc., produce the same resistance function within the immune system CRISPR.
@@ 2 Related Works
Research on Metagenomics. Traditional alignment-based methods like MetaPhlAn5 [57] aim to match similarities between query sequences and known reference genomes and are common for taxonomic profiling. Advancements in deep learning have led to new methods like CNN-MGP [2] and DeepVirFinder [54], which use CNNs for gene and viral classifications with one-hot encoding. K-mer tokenization [14], employed in approaches like MDL4Microbiome [34], is a standard for DNA sequence characterization. Additionally, Virtifier [46] maps a nucleotide sequence using a codon dictionary combined with LSTM to predict viral genes. DeepMicrobes [35] employs a self-attention mechanism, while DeepTE [62] uses K-mer inputs with CNNs for element classification, and Genomic-nlp [47] applies word2vec for gene function analysis. MetaTransformer [60] uses K-mer embedding for species classification with Transformer. For pre-training models, LookingGlass [20] uses a three-layer LSTM model for functional prediction in short DNA reads. ViBE [18] employs a K-mer token-based BERT model pre-trained with Masked Language Modeling for virus identification.
Pre-Training on Genomics. The BERT model, effective in DNA sequence characterization, is limited by the Transformer architecture’s computational burden. LOGO [64] addresses this by cutting off long sequences into 1-2kb sub-sequences. Enformer [6] combines extended convolution with Transformers for long human genomic data. GenSLMs [68] introduce hierarchical language models for whole-genome modelling. DNABERT [25], the first pre-trained model on the human genome that focuses on extracting efficient genomic representations. DNABERT2 [67], its successor, uses Byte Pair Encoding on multi-species genomic data. NT [10] is trained on nucleotide sequences from humans and other species and evaluated on 18 genome prediction tasks. HyenaDNA [48] presents a long-range genomic model based on single-nucleotide polymorphisms on human reference genomes.
@@ 3 Methods
In this section, we provide a detailed description of the proposed pre-training model FGBERT, which contains the MGM and TMC components as depicted in Fig. 2.
@@
Notation. Given a dataset of metagenomic long sequences $\{\mathcal{X}_{i}\}_{i=1}^{m}$, we simplify each $\mathcal{X}_{i}$ into a set of shorter gene sequences $\{{x_{i}}\}_{i=1}^{n_{i}}$ using ultrasonic fragmentation and assembly techniques [32], where $n_{i}$ represents the variable number of gene sequences derived from each $\mathcal{X}_{i}$. Each gene sequence $x_{i}$ is tokenized in a vector $g_{i}\in\mathbb{R}^{d}$, where $d$ is the token dimension. Each sequence is associated with a reduced-dimensional representation $y_{i}\in\mathbb{R}^{100}$. Suppose a gene group $\mathcal{G}={\left\{g_{i}\right\}}_{i=1}^{N}$ is formed by concatenating $N$ gene vectors sequentially. During the pre-training phase, each gene token $g_{i}\in\mathcal{G}$ is processed by the context-aware genome language encoder $\mathcal{F}(\cdot)$, generating the knowledge representations $z_{i}$, where $z_{i}=\mathcal{F}(g_{i})$. These representations are then transformed by a fully connected layer into $\hat{z}_{i}$, defined as $\hat{z}_{i}=\mathcal{H}(z_{i})$, where $\mathcal{H(\cdot)}$ represents a multi-classification head. In addition, we incorporate contrastive learning into the methodology. For each gene $x_{i}$, we introduce a data augmentation module to generate positive samples $x_{j(i)}$ and a hard negative sampling strategy for constructing negative samples $x_{k(i)}$.
3.1 Context-aware Masked Gene Modeling (MGM) for One-to-Many problem
Context-Aware Tokenizer. To develop a biologically meaningful tokenization of long sequences, we design a context-aware tokenizer utilizing the Protein Language Model (PLM), such as ESM-2 [36] with 15 billion parameters, integrating biological prior knowledge. As illustrated in Fig. 1, this tokenizer framework begins by extracting DNA gene sequences $\{{x_{i}}\}_{i=1}^{n_{i}}$ from a metagenomic sequence $\mathcal{X}_{i}$ utilizing the European Nucleotide Archive (ENA) software [17]. This conversion enhances the flexibility of analyzing longer metagenomic sequences.
Table 1: Framework of Context-Aware Tokenizer.
Secondly, these DNA sequences $x_{i}$ are translated into Amino Acid (AA) sequences using the Transeq software [45]. This translation addresses the issue of degenerate DNA codes, where certain non-‘ATCG’ symbols like ‘Y’ or ‘S’ can represent multiple nucleotides (e.g., ‘Y’ can be ‘C’ or ‘G’). By translating to AA sequences, we eliminate this redundancy, as different DNA sequences can map to the same AA sequence, ensuring consistency in representing biologically equivalent sequences [33, 24]. Thirdly, these AA sequences are transformed into 1280D normalized ESM-2 representations, with an additional gene orientation vector, resulting in 1281D gene representations $\{g_{i}\}_{i=1}^{n_{i}}$. The utility of ESM-2 lies in its ability to reveal gene relationships and functional information inherent in metagenomic data [53], preserving important intra-gene contextually-aware information. Finally, these representations every $N$ representations are concatenated sequentially to form gene groups $\mathcal{G}$, which serve as the basis for subsequent modeling tasks.
Masked Gene Modeling. We propose the MGM to enhance the model’s understanding of the relationships between genes within metagenomic sequences and their function regulations across diverse genomic environments (OTM problem). During pre-training, each gene token is masked with a 15% probability and predicted based on its unmasked genome context ${\mathcal{G}_{/M}}$:
$\mathcal{L}_{\text{MLM}}=\mathbb{E}_{g\sim\mathcal{G}}\mathbb{E}_{M}\sum_{i\in M}-\log p(g_{i}|{\mathcal{G}_{/M}}),$ (1)
where $M$ denotes the index set of the masked gene.
In addition, considering genetic polymorphism [52, 66], the MGM component focuses on detecting multiple genes that could coexist at a single genomic site, denoted as $\hat{z}_{i}=[\hat{z}_{1},\hat{z}_{2},\hat{z}_{3},\hat{z}_{4}]$. This requires the model not only to predict individual genes but also to identify various combinations of genes occurring within the same site. Thus, we enhance the model with a comprehensive loss function, $\mathcal{L}_{\text{MGM}}$, which incorporates feature reconstruction and probability prediction:
@@ 3.1 Context-aware Masked Gene Modeling (MGM) for One-to-Many problem
$\mathcal{L}_{\text{MGM}}=\frac{1}{N}\textstyle\sum\limits_{i=1}^{N}{(1-\frac{y_{i}^{T}\hat{z}_{i}}{\left\|y_{i}\right\|\cdot\left\|\hat{z}_{i}\right\|})}^{\gamma}+\\ \frac{\alpha}{N}\textstyle\sum\limits_{i=1}^{N}||\tilde{z}_{i}-\tilde{y}_{i}||_{2}^{2},$ (2)
@@
where $\gamma$ is a reconstruction loss with the scaled cosine error, and $\alpha$ is a weighting factor to balance the importance of the two loss functions. The first item, Feature Reconstruction Loss (FRL), quantifies the distance between the model prediction $\hat{z}_{i}$ and its corresponding label $y_{i}$. The second item, Probability Prediction Loss (PPL), evaluates the discrepancy between the predicted embedding probability $\tilde{z}_{i}=\frac{e^{\hat{z}_{i}}}{\sum_{j=1}^{C}e^{\hat{z}_{j}}}$ and the true category probability $\tilde{y}_{i}=\frac{e^{y_{i}}}{\sum_{j=1}^{C}e^{y_{j}}}$, both processed via the softmax function. $C$ denotes the number of gene combinations and is set to 4.
3.2 Triplet Metagenomic Contrastive Framework (TMC) for Many-to-One problem